OpenStudy (jojokiw3):

What is the equation of the circle that has its center at (0,0) and is tangent to the line with equation 3x-4y = 10?

OpenStudy (campbell_st):

and the equation of the circle is \[x^2 + y^2 = r^2\] the goal is to find r, the radius of the circle... so they there is an equation that lets you find the perpendicular distance from a line to a point. the tangent is always perpendicular to a radius... so if you find the distance from 3x - 4y = 10 to the point (0, 0) you'll be able to find r.... the formula is \[d = \frac{\left| Ax + By + C \right|}{\sqrt{A^2 + B^2}}\] so you have x = 0, y = 0, A = 3, B = -4 and C = -10 hope it helps